Find the equation of the tangent plane of the following surface patches at the indicated points: (a) σ(u,v) = (u,v,u2 + v2), P = (1,√2,3).
(b) σ(r,θ) = (rcosθ,rsinθ,2θ), P =(√3,1, π 3).
1
Expert's answer
2020-10-15T18:44:30-0400

(a) "\sigma (u,v)=(u,v,u^2+v^2)" , "P=(1,\sqrt{2}, 3)=\sigma(1,\sqrt{2})"

We have that "\sigma_u=(1,0,2u)" and "\sigma_v=(0,1,2v)" .

So, "\sigma_u(1,\sqrt{2})=(1,0,2)" and "\sigma _v(1,\sqrt{2})=(0,1,2\sqrt{2})"

Now we find the normal to the plane

"n=\sigma_u(1,\sqrt{2})\times \sigma_v(1,\sqrt{2})=\begin{vmatrix}i &j&k
\\ 1&0&2\\
0&1&2\sqrt{2}
\end{vmatrix}=(-2,-2\sqrt{2},1)"

The equation of the tangent plane at point "P" is:

"-2(x-1)-2\sqrt{2}(y-\sqrt{2})+1(z-3)=0"

or "2x+2\sqrt{2}y-z-3=0"


(b) "\sigma(r,\theta)=(r\cos \theta,r\sin \theta, 2\theta)" , "P(\sqrt{3}, 1,\pi/3)=\sigma(2,\pi/6)"

We have that "\sigma_r=(\cos\theta,\sin\theta, 0)" and "\sigma_\theta=(-r\sin\theta, r\cos\theta,2)"

So, "\sigma_r(2,\pi/6)=(\sqrt{3}/2,1/2,0)" and "\sigma_\theta(2,\pi/6)=(-1,\sqrt{3},2)"

Now we find the normal to the plane

"n=\sigma_r(2,\pi/6)\times \sigma_\theta(2,\pi/6)=\begin{vmatrix}i &j&k
\\ \sqrt{3}/2&1/2&0\\
-1&\sqrt{3}&2
\end{vmatrix}=(1,-\sqrt{3},2)"

The equation of the tangent plane at point "P" is:

"1(x-\sqrt{3})-\sqrt{3}(y-1)+2(z-\pi/3)=0"

or "x-\sqrt{3}y+2z-2\pi/3=0"


Answer: (a) "2x+2\sqrt{2}y-z-3=0"

(b) "x-\sqrt{3}y+2z-2\pi/3=0"